OpenStudy (anonymous):

Simplify: the square root of 2 (x − the square root of 5) i need help with these please

OpenStudy (anonymous):

Do you mean: \[\sqrt{2(x-5)}\]

OpenStudy (anonymous):

no ill write it on sec

OpenStudy (anonymous):

\[\sqrt{2x}-\sqrt{5}\]?

OpenStudy (anonymous):

\[\sqrt{2}(x-\sqrt{5})\]

OpenStudy (anonymous):

this one

OpenStudy (ranga):

Just distribute. Multiply each term inside parenthesis by sqrt(2). Simplify.

OpenStudy (anonymous):

\[\sqrt{2x}-\sqrt{10}\] ??????

OpenStudy (ranga):

x should be outside the root. And that's about it.

OpenStudy (anonymous):

\[x \sqrt{2}-\sqrt{10}\] ??/

OpenStudy (ranga):

Yeah. Or \[\Large \sqrt{2}x - \sqrt{10}\]

OpenStudy (anonymous):

but which one is the right one

OpenStudy (anonymous):

because both is on there its multiple choice

OpenStudy (ranga):

Either one is fine. But constant before the xterm is preferred.

OpenStudy (ranga):

Are you reading the choices correctly? I don't think both is likely to be there.

OpenStudy (anonymous):

yup both are on there i can take a screen shot if you like

OpenStudy (ranga):

I'd like to see that.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

see

OpenStudy (ranga):

Nope, it is not!

OpenStudy (anonymous):

yeah huhh

OpenStudy (ranga):

The last choice is the correct answer. The third choice has both 2 and x inside the square root. The one I showed you above has 2 inside the square root but x outside the square root.

OpenStudy (anonymous):

ok can you help me with mor eplease

OpenStudy (ranga):

one more. go ahead.

OpenStudy (anonymous):

Find the value of The square root of the quantity 8 over 9.

OpenStudy (anonymous):

one more?? :[

OpenStudy (ranga):

Use Equation editor so the problem will be clear.

OpenStudy (anonymous):

\[\sqrt{\frac{ 8 }{ 9 }}\]

OpenStudy (anonymous):

\[2\sqrt{\frac{ 3 }{ 25 }}\] i have to do the same thing to this

OpenStudy (ranga):

Rewrite numerator as 4 * 2 and the denominator as 3 * 3. See what all can be taken out of the square root.

OpenStudy (anonymous):

huh??

OpenStudy (anonymous):

would it be \[\frac{ 2 }{ 3 }\]

OpenStudy (ranga):

\[\sqrt{\frac{ 8 }{ 9 }} = \frac{ \sqrt{4 * 2} }{ \sqrt{9} } = \frac{ \sqrt{4} * \sqrt{2} }{ \sqrt{9} }\]

OpenStudy (ranga):

It will be 2/3 * sqrt(2)

OpenStudy (ranga):

\[\frac{ 2\sqrt{2} }{ 3 }\]

OpenStudy (anonymous):

ohh ok thank you

OpenStudy (ranga):

Welcome.

OpenStudy (anonymous):

i might need helo a lil bit later on

OpenStudy (ranga):

If I am around I will try to help.

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

@ranga i need you

OpenStudy (ranga):

What is the question?

OpenStudy (anonymous):

Find the value of \[\frac{ 3x }{ \sqrt{x^{3}} }\] Find the value of \[\frac{ 2\sqrt{3} }{ \sqrt{10} }\] in simplest form

OpenStudy (ranga):

For the first one, write denominator as sqrt( x^2 * x ). The reason we do that is because we can easily take the square root of a perfect square. When you take sqrt(x^2) you will end up with x. So the denominator will be x * sqrt(x). Write that step first. Then I will tell you the next step.

OpenStudy (anonymous):

3 over sqrt x

OpenStudy (ranga):

Correct. The next step is to "rationalize the denominator." That means not to leave any radical sign such as square root, cube root, etc. in the denominator. Multiply both numerator and denominator by sqrt(x) to get rid of the radical sign in the denominator. What do you get?

OpenStudy (anonymous):

wait is this for the second one

OpenStudy (anonymous):

3 over x

OpenStudy (ranga):

No the first one. They are different forms of answers. The preferred form is not to leave radical in the denominator. But sometimes they accept that answer but a lot of times they don't.

OpenStudy (ranga):

You have to multiply BOTH the numerator and denominator by sqrt(x). Looks like you only multiplied the denominator to get 3/x.

OpenStudy (anonymous):

ohhh \[\frac{ 3\sqrt{x} }{ x }\]

OpenStudy (ranga):

Is that one of your answer choices?

OpenStudy (anonymous):

yes

OpenStudy (ranga):

That is the correct answer.

OpenStudy (ranga):

Now you should be able to do the other problem knowing that we should not have radicals in the denominator. What must you do to get rid of the radical in denominator?

OpenStudy (anonymous):

multiply both the numerator and denominator?

OpenStudy (ranga):

yes, by what?

OpenStudy (anonymous):

the radical?

OpenStudy (ranga):

by square root of 10.

OpenStudy (anonymous):

yeah so it becomes \[\frac{ 2\sqrt{30} }{ 10 }\]

OpenStudy (ranga):

Yes. 2/10 can be simplified further.

OpenStudy (anonymous):

it can?

OpenStudy (anonymous):

2/5?

OpenStudy (ranga):

come on. try again.

OpenStudy (anonymous):

1/5

OpenStudy (ranga):

yes. check the answer after simplification.

OpenStudy (anonymous):

\[\frac{ \sqrt{30} }{ 5 }\]

OpenStudy (ranga):

Yes. Is it one of the answer choices?

OpenStudy (anonymous):

yes

OpenStudy (ranga):

That is it. ttyl.

OpenStudy (anonymous):

ok bye thank you

OpenStudy (ranga):

you are welcome.